# Transparent Image Layer Diffusion using Latent Transparency

### v3

@@
Abstract
We present LayerDiffuse, an approach enabling large-scale pretrained latent diffusion models to generate transparent images. The method allows generation of single transparent images or of multiple transparent layers. The method learns a “latent transparency” that encodes alpha channel transparency into the latent manifold of a pretrained latent diffusion model. It preserves the production-ready quality of the large diffusion model by regulating the added transparency as a latent offset with minimal changes to the original latent distribution of the pretrained model. In this way, any latent diffusion model can be converted into a transparent image generator by finetuning it with the adjusted latent space. We train the model with 1M transparent image layer pairs collected using a human-in-the-loop collection scheme. We show that latent transparency can be applied to different open source image generators, or be adapted to various conditional control systems to achieve applications like foreground/background-conditioned layer generation, joint layer generation, structural control of layer contents, etc. A user study finds that in most cases (97%) users prefer our natively generated transparent content over previous ad-hoc solutions such as generating and then matting. Users also report the quality of our generated transparent images is comparable to real commercial transparent assets like Adobe Stock.
1 Introduction
”Woman with messy hair, in the bedroom”
”Burning firewood, on a table, in the countryside”
Blended output
Output layer 1
Output layer 2
Figure 1: Generating transparent images and layers. For the given text prompts (top), our framework is capable of generating multiple layers with transparency. These layers can be blended to produce images corresponding to the prompts. Zoom in to see details including messy hair and semi-transparent fire.
@@ 1 Introduction
The primary factors contributing to this research gap are the lack of training data and the difficulty in manipulating the data representation of existing large-scale image generators. High-quality transparent image elements on the Internet are typically hosted by commercial image stocks with limited (and costly) access, in contrast to text-image datasets that already include billions of images (e.g., LAION [39]). The largest open-source transparent image datasets are often less than 50K in size (e.g., DIM [54] includes 45,500 transparent images). Meanwhile, most open-source image generation models, e.g., Stable Diffusion, are latent diffusion models that are sensitive to their latent space data representations. Even minor changes to the latent distribution could severely degrade inference or finetuning. For instance, Stable Diffusion 1.5 and XL use different latent spaces, and finetuning with mismatched latents can cause significant degradation in output image quality [45]. This adds to the challenge of manipulating the data representation of existing models to support additional formats like transparent images.
We present a ”latent transparency” approach that enables large-scale pretrained latent diffusion models to generate transparent images as well as multiple transparent layers. This method encodes image transparency into a latent offset that is explicitly regulated to avoid disrupting the latent distribution. The latent transparency is encoded and decoded by external independent models, ensuring that the original pretrained latent encoder/decoder is perfectly preserved, so as to maintain high-quality results of state-of-the-art diffusion models. To generate multiple layers together, we use a shared attention mechanism to ensure consistency and harmonious blending between image layers, and train LoRAs to adapt the models to different layer conditions.
@@
Experiments show that in a dominant majority of cases (97%), users prefer the transparent content generated natively by our method over previous ad-hoc solutions like generating-then-matting. When we compare the quality of our generated results with the search results from commercial transparent assets sites like Adobe Stock, user preference rates suggest that quality is comparable.
@@
In summary, we (1) propose “latent transparency”, an approach to enable large-scale pretrained latent diffusion models to generate single transparent images or multiple transparent layers, (2) present a shared attention mechanism to generate layers with consistent and harmonious blending, (3) present a dataset containing transparent images and layers collected using a human-in-the-loop scheme, (4) present a pretrained model for transparent image generation and two pretrained LoRAs for multiple layer generation.
2 Related Work
2.1 Hiding Images inside Perturbations
@@
Research in multiple fields point out a phenomenon: neural networks have the ability to “hide” features in perturbations inside existing features without changing the overall feature distributions, e.g., hiding an image inside another image through small, invisible pixel perturbations. A typical CycleGAN [62] experiment showcases face-to-ramen, where the human face identity could be hidden in a ramen picture. Similarly, invertible downscaling [53] and invertible grayscale [52] indicate that neural networks can hide a large image inside a smaller one, or hide a colorful image inside a grayscale one, and then reconstruct the original image. In another widely verified experiment Goodfellow et al. [14] show that adversarial example signals can be hidden inside feature perturbations to influence the behaviors of other neural networks. In this paper, our proposed “latent transparency” utilizes similar principles: hiding image transparency features inside a small perturbation added to the latent space of Stable Diffusion [44], while at the same time avoiding changes to the overall distribution of the latent space. This enables a seamless transition from existing non-transparent image diffusion models to transparent diffusion models, and eases further fine-tuning.
2.2 Diffusion Probabilistic Models and Latent Diffusion
@@
Diffusion Probabilistic Model [41] and related training and sampling methods like Denoising Diffusion Probabilistic Model (DDPM) [16], Denoising Diffusion Implicit Model (DDIM) [42], and score-based diffusion [43] contribute to the foundations of recent large-scale image generators. Early image diffusion methods usually directly use pixel colors as training data [42, 37, 24]. Recently Latent Diffusion Model (LDM) [35] has been proposed for easier training and lowering the computation requirements. This method has been further extended to create Stable Diffusion [44]. Even more recently, eDiff-I [6] uses an ensemble of multiple conditions including a T5 text encoder [33], a CLIP text and image embedding encoder [19]. Versatile Diffusion [55] adopts a multi-purpose diffusion framework to process text, an image, and variations within a single model. To achieve richer functionality, Composer [18] uses a joint training method with various control signals over the pretrained image diffusion model conditioned on image embedding. Mixture-of-Experts (MoEs) methods [40, 12] also contribute to the models of text-to-image generation.
2.3 Customized Diffusion Models and Image Editing
Early methods to customize diffusion models have focused on text-guidance [31, 21, 5, 7, 20, 34, 15]. Image diffusion algorithms also naturally support inpainting [34, 5]. Textual Inversion [13] and DreamBooth [36] can personalize the contents of generated results based on a small set of examplar images of the same topic or object. Recently, control models have also been used to add additional conditions for the generation of text-to-image models, e.g., ControlNet [59], lightweight T2I-adapter [30], etc. Uni-ControlNet [61] proposes a multi-scale condition injection strategy to learn an adapter for various local controls to reduce the training computation costs. SeeCoder [56] presents a semantic context encoder to replace the original text encoder to generate image variants. IP-Adapter[58] uses a cross-attention mechanism to separate text and image features, allowing for the control signals of the reference image as a visual prompt. More recently researchers  [51, 28] have proposed portrait personalization methods using one or more face images. [60] attempts to use image segmentation and inpainting to achieved layered effects in diffusion models. [27] uses masks in neural network features to achieve semantic region control. Experiments show that our approach can be combined with various control models or customization methods to generate transparent layers for different applications.
2.4 Transparent Image Processing
Transparent image processing is closely related to image decomposition, layer extraction, color palette processing, as well as image matting [50, 2, 1]. Typical color-based decomposition can be viewed as a RGB color space geometry problem [46, 47, 48, 49, 11]. These ideas have also been extended to more advanced blending of image layers [25]. Unmixing-based color separation also contributes to image decomposition [3], and semantic features can be used in image soft segmentation [4]. We compare our approach to several state-of-the-art deep-learning based matting methods in our experiments and discussion. PPMatting [9] is a neural network image matting model trained from scratch using standard matting datasets. Matting Anything [26] is a image matting model using the Segment Anything Model (SAM) [22] as a backbone. VitMatte [57] is a trim-map-based matting method using Vision Transformer (ViT).
@@ 2.4 Transparent Image Processing
Figure 2: Latent Transparency. Given an input transparent image, our framework encode a “latent transparency” to adjust the latent space of Stable Diffusion. The adjusted latent images can be decoded to reconstruct the color and alpha. This latent space with transparency can be further used in training or fine-tuning pretrained image diffusion models.
3 Method
@@
Our approach enables a Latent Diffusion Model (LDM), like Stable Diffusion, to generate transparent images, and then further extend the model to jointly generate multiple transparent layers together. In section 3.1, we introduce the method to adjust the LDM latent space to support transparent image encoding/decoding. In section 3.2, we adapt pretrained latent diffusion models with the adjusted latent space to generate transparent images. In section 3.3, we describe the method for joint or conditional layer generating. Finally, we detail the dataset preparation and implementation details of neural network training in section 3.4.
@@
To clarify the presentation we first define some terms. For any transparent image $\bm{I}_{t}\in\mathbb{R}^{h\times w\times 4}$ with RGBA channels, we denote the first 3 RGB color channels as $\bm{I}_{c}\in\mathbb{R}^{h\times w\times 3}$ and the alpha channel as $\bm{I}_{\alpha}\in\mathbb{R}^{h\times w\times 1}$. Since the colors are physically undefined at pixels where the alpha value is strictly zero, in this paper, all undefined areas in $\bm{I}_{c}$ are always padded by an iterative Gaussian filter (Appendix A) to avoid aliasing and unnecessary edge patterns. We call $\bm{I}_{c}$ the “padded RGB image” (Fig. 2). The $\bm{I}_{t}$ can be converted to a “premultiplied image” as $\bm{I}=\bm{I}_{c}*\bm{I}_{a}$ where *** denotes pixelwise multiplication. In this paper, all RGB values are in range $[-1,1]$ (consistent with Stable Diffusion) while all alpha values are in range $[0,1]$. The premultiplied image $\bm{I}$ can be seen as a common non-transparent RGB image that can be processed by any RGB-formatted neural networks. Visualizations of these images are shown in Fig. 2.
3.1 Latent Transparency
Our goal is to add transparency support to large-scale latent diffusion models, like Stable Diffusion (SD), that typically uses a latent encoder (VAE) to convert RGB images to latent images before feeding it to a diffusion model. Herein, the VAE and the diffusion model should share the same latent distribution, as any major mismatch can significantly degrade the inference/training/fine-tuning of the latent diffusion framework. When we adjust the latent space to support transparency, the original latent distribution must be preserved as much as possible. This seemingly ill-defined objective can be determined by a straight-forward measurement: we can check how well the modified latent distribution can be decoded by the original pretrained frozen latent decoder — if decoding a modified latent image creates severe artifacts, the latent distribution is misaligned or broken.
We can write this ”harmfulness” measurement mathematically as follows. Given an RGB image $\bm{I}$, the pretrained and frozen Stable Diffusion latent encoder $\mathcal{E}^{*}_{sd}(\cdot)$ and decoder $\mathcal{D}^{*}_{sd}(\cdot)$, where the *** indicates frozen models, we denote the latent image as $\bm{x}=\mathcal{E}^{*}_{sd}(\bm{I})$. Assuming this latent image $\bm{x}$ is modified by any offset $\bm{x}_{\epsilon}$, produces an adjusted latent $\bm{x}_{a}=\bm{x}+\bm{x}_{\epsilon}$. The decoded RGB reconstruction can then be written as $\hat{\bm{I}}=\mathcal{D}^{*}_{sd}(\bm{x}_{a})$ and we can evaluate how “harmful” the offset $\bm{x}_{\epsilon}$ is as
$\mathcal{L}_{\text{identity}}=||\bm{I}-\hat{\bm{I}}||_{2}=||\bm{I}-\mathcal{D}^{*}_{sd}(\mathcal{E}^{*}_{sd}(\bm{I})+\bm{x}_{\epsilon})||_{2}\,,$ (1)
@@ 3.1 Latent Transparency
$\bm{x}_{\epsilon}=\mathcal{E}(\bm{I}_{c},\bm{I}_{\alpha})\,.$ (2)
@@
$[\hat{\bm{I}_{c}}\,\,\,\hat{\bm{I}_{\alpha}}]=\mathcal{D}(\hat{\bm{I}},\bm{x}_{a})\,,$ (3)
where $\hat{\bm{I}_{c}},\hat{\bm{I}_{\alpha}}$ are the reconstructed color and alpha channels. The neural network layer architecture of $\mathcal{E}(\cdot,\cdot)$ and $\mathcal{D}(\cdot,\cdot)$ is in Appendix B. We evaluate the reconstruction with
$\mathcal{L}_{\text{recon}}=||\bm{I_{c}}-\hat{\bm{I}_{c}}||_{2}+||\bm{I_{a}}-\hat{\bm{I}_{a}}||_{2}\,,$ (4)
@@
$\mathcal{L}_{\text{disc}}=\mathbb{L}_{\text{disc}}([\hat{\bm{I}_{c}},\hat{\bm{I}_{a}}])\,,$ (5)
where $\mathbb{L}_{\text{disc}}(\cdot,\cdot)$ is a GAN objective from a 5-layer patch discriminator (Appendix C). The final objective can be jointly written as
$\mathcal{L}_{\text{vae}}=\lambda_{\text{recon}}\mathcal{L}_{\text{recon}}+\lambda_{\text{identity}}\mathcal{L}_{\text{identity}}+\lambda_{\text{disc}}\mathcal{L}_{\text{disc}}\,,$ (6)
@@
3.2 Diffusion Model with Latent Transparency
Since the altered latent space with latent transparency is explicitly regulated to align with the original pretrained latent distribution (Eq. 1), Stable Diffusion can be directly fine-tuned on the altered latent space. Given the adjusted latent $\bm{x}_{a}$, diffusion algorithms progressively add noise to the image and produces a noisy image $\bm{x}_{t}$, with $t$ denoting how many times noise is added. When $t$ is large enough, the latent image approximates pure noise. Given a set of conditions including the time step $t$ and text prompts $\bm{c}_{t}$, image diffusion algorithms learn a network $\epsilon_{\theta}$ that predicts the noise added to the noisy latent image $\bm{x}_{t}$ with
$\mathcal{L}=\mathbb{E}_{\bm{x}_{t},t,\bm{c}_{t},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{x}_{t},t,\bm{c}_{t}))\|_{2}^{2}\Big{]}$ (7)

### v4

@@
Abstract.
We present an approach enabling large-scale pretrained latent diffusion models to generate transparent images. The method allows generation of single transparent images or of multiple transparent layers. The method learns a “latent transparency” that encodes alpha channel transparency into the latent manifold of a pretrained latent diffusion model. It preserves the production-ready quality of the large diffusion model by regulating the added transparency as a latent offset with minimal changes to the original latent distribution of the pretrained model. In this way, any latent diffusion model can be converted into a transparent image generator by finetuning it with the adjusted latent space. We train the model with 1M transparent image layer pairs collected using a human-in-the-loop collection scheme. We show that latent transparency can be applied to different open source image generators, or be adapted to various conditional control systems to achieve applications like foreground/background-conditioned layer generation, joint layer generation, structural control of layer contents, etc. A user study finds that in most cases (97%) users prefer our natively generated transparent content over previous ad-hoc solutions such as generating and then matting. Users also report the quality of our generated transparent images is comparable to real commercial transparent assets like Adobe Stock.
Transparent images, image editing, image layer, text-to-image diffusion
submissionid: 279journal: TOGcopyright: acmlicensedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 4article: 100publicationmonth: 7doi: 10.1145/3658150ccs: Applied computing Fine artsccs: Applied computing Media arts
”Woman with messy hair, in the bedroom” ”Burning firewood, on a table, in the countryside”
Blended output Output layer 1 Output layer 2 Blended output Output layer 1 Output layer 2
Figure 1. Generating transparent images and layers. For the given text prompts (top), our framework is capable of generating multiple layers with transparency. These layers can be blended to produce images corresponding to the prompts. Zoom in to see details including messy hair and semi-transparent fire.
1. Introduction
@@
The primary factors contributing to this research gap are the lack of training data and the difficulty in manipulating the data representation of existing large-scale image generators. High-quality transparent image elements on the Internet are typically hosted by commercial image stocks with limited (and costly) access, in contrast to text-image datasets that already include billions of images (e.g., LAION (Schuhmann et al., 2022)). The largest open-source transparent image datasets are often less than 50K in size (e.g., DIM (Xu et al., 2017) includes 45,500 transparent images). Meanwhile, most open-source image generation models, e.g., Stable Diffusion, are latent diffusion models that are sensitive to their latent space data representations. Even minor changes to the latent distribution could severely degrade inference or finetuning. For instance, Stable Diffusion 1.5 and XL use different latent spaces, and finetuning with mismatched latents can cause significant degradation in output image quality (Stability, 2022b). This adds to the challenge of manipulating the data representation of existing models to support additional formats like transparent images.
We present a ”latent transparency” approach that enables large-scale pretrained latent diffusion models to generate transparent images as well as multiple transparent layers. This method encodes image transparency into a latent offset that is explicitly regulated to avoid disrupting the latent distribution. The latent transparency is encoded and decoded by external independent models, ensuring that the original pretrained latent encoder/decoder is preserved, so as to maintain high-quality results of state-of-the-art diffusion models. To generate multiple layers together, we use a shared attention mechanism that ensures consistency and harmonious blending between image layers, and we train LoRAs to adapt the models to different layer conditions.
@@
Experiments show that in a majority of cases (97%), users prefer the transparent content generated natively by our method over previous ad-hoc solutions like generating-then-matting. When we compare the quality of our generated results with the search results from commercial transparent assets sites like Adobe Stock, user preference rates suggest that quality is comparable.
@@
In summary, we (1) propose “latent transparency”, an approach to enable large-scale pretrained latent diffusion models to generate single transparent images or multiple transparent layers, (2) we present a shared attention mechanism to generate layers with consistent and harmonious blending, and (3) we present a pretrained model for transparent image generation, two pretrained LoRAs for multiple layer generation, as well as several additional ablative architectures for multi-layer generation.
2. Related Work
2.1. Hiding Images inside Perturbations
@@
Research in multiple fields point out a phenomenon: neural networks have the ability to “hide” features in perturbations inside existing features without changing the overall feature distributions, e.g., hiding an image inside another image through small, invisible pixel perturbations. A typical CycleGAN (Zhu et al., 2017) experiment showcases face-to-ramen, where the human face identity could be hidden in a picture of ramen. Similarly, invertible downscaling (Xiao et al., 2020) and invertible grayscale (Xia et al., 2018) indicate that neural networks can hide a large image inside a smaller one, or hide a colorful image inside a grayscale one, and then reconstruct the original image. In another widely verified experiment Goodfellow et al. (2015) show that adversarial example signals can be hidden inside feature perturbations to influence the behaviors of other neural networks. In this paper, our proposed “latent transparency” utilizes similar principles: hiding image transparency features inside a small perturbation added to the latent space of Stable Diffusion (Stability, 2022a), while at the same time avoiding changes to the overall distribution of the latent space.
2.2. Diffusion Probabilistic Models and Latent Diffusion
@@
Diffusion Probabilistic Model (Sohl-Dickstein et al., 2015) and related training and sampling methods like Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020), Denoising Diffusion Implicit Model (DDIM)  (Song et al., 2021), and score-based diffusion  (Song et al., 2020) contribute to the foundations of recent large-scale image generators. Early image diffusion methods usually directly use pixel colors as training data (Song et al., 2021; San-Roman et al., 2021; Kong and Ping, 2021). In contrast, the Latent Diffusion Model (LDM) (Rombach et al., 2022) operates in latent space and has been shown to enable easier training while lowering computation requirements. This method has been further extended to create Stable Diffusion (Stability, 2022a). Recently, eDiff-I (Balaji et al., 2022) has used an ensemble of multiple conditions including a T5 text encoder (Raffel et al., 2019), a CLIP text and image embedding encoder (Ilharco et al., 2021). Versatile Diffusion (Xu et al., 2022) adopts a multi-purpose diffusion framework to process text, an image, and variations within a single model.
2.3. Customized Diffusion Models and Image Editing
Early methods to customize diffusion models have focused on text-guidance (Nichol et al., 2021; Kim et al., 2022a; Avrahami et al., 2022). Image diffusion algorithms also naturally support inpainting (Ramesh et al., 2022; Avrahami et al., 2022). Textual Inversion (Gal et al., 2022) and DreamBooth (Ruiz et al., 2022) can personalize the contents of generated results based on a small set of examplar images of the same topic or object. Recently, control models have also been used to add additional conditions for the generation of text-to-image models, e.g., ControlNet (Zhang and Agrawala, 2023), lightweight T2I-adapter (Mou et al., 2023), etc. IP-Adapter(Ye et al., 2023) uses a cross-attention mechanism to separate text and image features, allowing for the control signals of the reference image as a visual prompt. (Li et al., 2023a) uses masks in neural network features to achieve semantic region control. Inversion-based methods are also popular in editing images. The DDPM (Ho et al., 2020) theory indicates that a diffusion algorithm constructs data with accumulated small variations and those variations, conditioned on noise, can be manipulated with inverted optimization. Mokady et al. (2023) shows that DDIM inversion can optimize images without requiring inputs to be generated by a previously known diffusion process (null-text embedding). Cao et al. (2023) and Narek et al. (2023) manipulate spatial cross-attention features of Stable Diffusion layers together with DDPM inversion. Hertz et al. (2023) edit attention activations of the input images with user-given text prompts and feed them back to the diffusion models. DiffEdit (2023) generates region masks for image editing, given input images and user prompts. DiffusionCLIP (Kim et al., 2022b) finetunes diffusion models with CLIP loss against prompts. Imagic (Kawar et al., 2023) jointly optimizes text embedding of user prompts and the model gradients to reconstruct the image for image editing applications.
@@ 2.3. Customized Diffusion Models and Image Editing
Figure 2. Latent Transparency. Given an input transparent image, our framework encode a “latent transparency” to adjust the latent space of Stable Diffusion. The adjusted latent images can be decoded to reconstruct the color and alpha. This latent space with transparency can be further used in training or fine-tuning pretrained image diffusion models.
2.4. Transparent Image Layer Processing
Transparent image processing is closely related to image decomposition, layer extraction, color palette processing, as well as image matting (Tang et al., 2019; Aksoy et al., 2017a, 2016). Typical color-based decomposition can be viewed as a RGB color space geometry problem (Tan et al., 2015, 2016, 2018, 2019; Du et al., 2023). These ideas have also been extended to more advanced blending of image layers (Koyama and Goto, 2018). Unmixing-based color separation also contributes to image decomposition (Aksoy et al., 2017b), and semantic features can be used in image soft segmentation (Aksoy et al., 2018). We compare our approach to several state-of-the-art deep-learning based matting methods in our experiments and discussion. PPMatting (Chen et al., 2022) is a neural network image matting model trained from scratch using standard matting datasets. Matting Anything (Li et al., 2023b) is a image matting model using the Segment Anything Model (SAM) (Kirillov et al., 2023) as a backbone. VitMatte (Yao et al., 2024) is a tri-map-based matting method using a Vision Transformer (ViT). Text2Layer (Zhang et al., 2023) attempts to use foreground segmentation guidance to achieved layered effects in diffusion models, and indicates that its main bottleneck is the quality of foreground matting method since its learning objective is constructed from the image segmentation of matting models. Our approach starts from native generation of transparent images rather than post-processing of image matting, and is fundamentally different from previous approaches that use matting as post-processing of model outputs or use matting for dataset synthesizing.
2.5. Image Harmonization
Harmonious blending of transparent image layers is closely related to image harmonization research. Achieving “harmony” is usually seen as a problem of correlating color, contrast, and style constituents between foreground and background to ensure natural appearance and consistent composition. Deep learning approaches (Zhu et al., 2015; Tsai et al., 2017; Guo et al., 2021; Chen et al., 2023; Guerreiro et al., 2023; Tan et al., 2023) have been proposed to harmonize images, using annotated datasets (Cong et al., 2020; Niu et al., 2023). These works utilize the learning capabilities of neural networks to acquire the prior knowledge of harmonization.
3. Method
@@
Our approach enables a Latent Diffusion Model (LDM), like Stable Diffusion, to generate transparent images, and then extends the model to jointly generate multiple transparent layers together. In section 3.1, we introduce the method to adjust the LDM latent space to support transparent image encoding/decoding. In section 3.2, we adapt pretrained latent diffusion models with the adjusted latent space to generate transparent images. In section 3.3, we describe the method for joint or conditional layer generating. Finally, we detail the dataset preparation and implementation details for neural network training in section 3.4.
@@
To clarify the presentation we first define some terms. For any transparent image $\bm{I}_{t}\in\mathbb{R}^{h\times w\times 4}$ with RGBA channels, we denote the first 3 RGB color channels as $\bm{I}_{c}\in\mathbb{R}^{h\times w\times 3}$ and the alpha channel as $\bm{I}_{\alpha}\in\mathbb{R}^{h\times w\times 1}$. Since the colors are physically undefined at pixels where the alpha value is strictly zero, in this paper, all undefined areas in $\bm{I}_{c}$ are always padded by an iterative Gaussian filter (see also supplementary material) to avoid aliasing and unnecessary edge patterns. We call $\bm{I}_{c}$ the “padded RGB image” (Fig. 2). The $\bm{I}_{t}$ can be converted to a “premultiplied image” as $\bm{I}=\bm{I}_{c}*\bm{I}_{a}$ where * denotes pixelwise multiplication. In this paper, all RGB values are in range $[-1,1]$ (consistent with Stable Diffusion) while all alpha values are in range $[0,1]$. The premultiplied image $\bm{I}$ can be seen as a common non-transparent RGB image that can be processed by any RGB-formatted neural networks. Visualizations of these images are shown in Fig. 2.
3.1. Latent Transparency
Our goal is to add transparency support to large-scale latent diffusion models, like Stable Diffusion (SD), that typically uses a latent encoder (VAE) to convert RGB images to latent images before feeding it to a diffusion model. Herein, the VAE and the diffusion model should share the same latent distribution, as any major mismatch can significantly degrade the inference/training/fine-tuning of the latent diffusion framework. When we adjust the latent space to support transparency, the original latent distribution must be preserved as much as possible. These seemingly conflicting goals (adding transparency support while preserving the original latent distribution) can be handled with a straight-forward measurement: we can check how well the modified latent distribution can be decoded by the original pretrained frozen latent decoder — if decoding a modified latent image creates severe artifacts, the latent distribution is misaligned or broken.
We can write this ”harmfulness” measurement mathematically as follows. Given an RGB image $\bm{I}$, the pretrained and frozen Stable Diffusion latent encoder $\mathcal{E}^{*}_{sd}(\cdot)$ and decoder $\mathcal{D}^{*}_{sd}(\cdot)$, where the * indicates frozen models, we denote the latent image as $\bm{x}=\mathcal{E}^{*}_{sd}(\bm{I})$. Assuming this latent image $\bm{x}$ is modified by any offset $\bm{x}_{\epsilon}$, produces an adjusted latent $\bm{x}_{a}=\bm{x}+\bm{x}_{\epsilon}$. The decoded RGB reconstruction can then be written as $\hat{\bm{I}}=\mathcal{D}^{*}_{sd}(\bm{x}_{a})$ and we can evaluate how “harmful” the offset $\bm{x}_{\epsilon}$ is as
(1) $\mathcal{L}_{\text{identity}}=||\bm{I}-\hat{\bm{I}}||_{2}=||\bm{I}-\mathcal{D}^{*}_{sd}(\mathcal{E}^{*}_{sd}(\bm{I})+\bm{x}_{\epsilon})||_{2}\,,$
@@ 3.1. Latent Transparency
Besides, since most mainstream VAE for diffusion models are KL-Divergence or Diagonal Gaussian Distribution models, these models often has a naively trained parameter for the standard deviation as an offset in the latent space. Considering such deviation denoted as $\bm{x}_{\text{std}}$, we can make use of this pretrained parameter to construct $\bm{x}_{\epsilon}=\lambda_{\text{offset}}\bm{x}_{\text{std}}\bm{x}_{\text{offset}}$ where $\bm{x}_{\text{offset}}$ is the raw output from newly added encoder, $\bm{x}_{\text{std}}$ is the deviation output of pretrained VAE, and $\lambda_{\text{offset}}$ is a weighting parameter with default $\lambda_{\text{offset}}=1e2$.
@@
(2) $\bm{x}_{\epsilon}=\mathcal{E}(\bm{I}_{c},\bm{I}_{\alpha})\,.$
@@
(3) $[\hat{\bm{I}_{c}}\,\,\,\hat{\bm{I}_{\alpha}}]=\mathcal{D}(\hat{\bm{I}},\bm{x}_{a})\,,$
where $\hat{\bm{I}_{c}},\hat{\bm{I}_{\alpha}}$ are the reconstructed color and alpha channels. The neural network layer architecture of $\mathcal{E}(\cdot,\cdot)$ and $\mathcal{D}(\cdot,\cdot)$ is in the supplementary material. We evaluate the reconstruction with
(4) $\mathcal{L}_{\text{recon}}=||\bm{I_{c}}-\hat{\bm{I}_{c}}||_{2}+||\bm{I_{a}}-\hat{\bm{I}_{a}}||_{2}\,,$
@@
(5) $\mathcal{L}_{\text{disc}}=\mathbb{L}_{\text{disc}}([\hat{\bm{I}_{c}},\hat{\bm{I}_{a}}])\,,$
where $\mathbb{L}_{\text{disc}}(\cdot,\cdot)$ is a GAN objective from a 5-layer patch discriminator (details in supplementary material). The final objective can be jointly written as
(6) $\mathcal{L}_{\text{vae}}=\lambda_{\text{recon}}\mathcal{L}_{\text{recon}}+\lambda_{\text{identity}}\mathcal{L}_{\text{identity}}+\lambda_{\text{disc}}\mathcal{L}_{\text{disc}}\,,$
@@
3.2. Diffusion Model with Latent Transparency
Figure 3. Model Training. We visualize the training of the base model to generate transparent images, and the training of the multi-layer model to generate multiple layers together. When training the base diffusion model (a), all model weights are trainable, whereas for training the multi-layer model (b), only two LoRAs are trainable (the foreground LoRA and background LoRA).
Since the altered latent space with latent transparency is explicitly regulated to align with the original pretrained latent distribution (Eq. 1), Stable Diffusion can be directly fine-tuned on the altered latent space. Given the adjusted latent $\bm{x}_{a}$, diffusion algorithms progressively add noise to the image and produce a noisy image $\bm{x}_{t}$, with $t$ denoting how many times noise is added. When $t$ is large enough, the latent image approximates pure noise. Given a set of conditions including the time step $t$ and text prompt $\bm{c}_{t}$, image diffusion algorithms learn a network $\epsilon_{\theta}$ that predicts the noise added to the noisy latent image $\bm{x}_{t}$ with
(7) $\mathcal{L}=\mathbb{E}_{\bm{x}_{t},t,\bm{c}_{t},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{x}_{t},t,\bm{c}_{t}))\|_{2}^{2}\Big{]}$